OpenStudy (anonymous):

Which one of the following sets of numbers represents the lengths of the sides of a right triangle? A. 13, 35, 37 B. 5, 10, 15 C. 9, 40, 41 D. 6, 9, 10 Can you tell me the reason too. I'll give metals.

jigglypuff314 (jigglypuff314):

Pythagorean theorem: a^2 + b^2 = c^2 so for example, does 6^2 + 9^2 = 10^2 ? 36 + 81 = 100 117 = 100 but 117 does not equal 100 therefore Option D, 6 9 10 will not work for a right triangle. then you can try plugging in like I had for the rest of them :)

OpenStudy (anonymous):

Thank you ! The answer is C right ?

jigglypuff314 (jigglypuff314):

Correct :)